OpenStudy (anonymous):

Solve ln (x+1)=ln x+1 for x

OpenStudy (tkhunny):

This might make it look easier: \(ln(x+1) = ln(x) + ln(e) \)

OpenStudy (dumbcow):

move logs to one side combine logs \[\ln (x+1) -\ln x = 1\] \[\ln \frac{x+1}{x} = 1\] take "e" raised to both sides or change to exponential form \[\frac{x+1}{x} = e\] \[x+1 = ex\] \[1 = ex -x\] \[x = \frac{1}{e-1}\]

OpenStudy (dumbcow):

@tkhunny , that would save a couple of steps, i didn't think of that

OpenStudy (anonymous):

Thank you both

OpenStudy (tkhunny):

I am always delighted when unique answers don't care how you find them. ALWAYS keep your mind open for alternatives.